OpenStudy (anonymous):

Solve and show work 3x − 8 = 2x + 12

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

3x-8=2x+12 -2x -2x 1x-8=12 -8 -8 1x=4 X=4

OpenStudy (anonymous):

oh my bad i meant to add the 8. you're right it is 20

OpenStudy (anonymous):

3x-8=2x+12 -2x -2x 1x-8=12 +8 +8 1x=20 X=20

OpenStudy (anonymous):

3x − 8 = 2x + 120 3x - 2x = 12 + 8 ->x = 20